Neutron stars are some of the densest objects in the universe. They form during supernova explosions and are very small compared to other astronomical objects. You can assume 5 x 1017 kg/m3 as the average density of a neutron star. The Earth has a total mass of about 5.97 x 1024 kg. What would be the diameter of the Earth if it had the density of a neutron star?


1
Expert's answer
2022-03-29T12:27:37-0400

"d=\sqrt[3]{\frac{6m}{\pi d}}=0.28~km."